if cose theta – cot theta =p then cosec + cot theta 

Grade:12th pass

2 Answers

Saurabh Koranglekar
4 years ago
Dear student

Cosec(x) sq - cot ( x) sq = 1

By identity

So answer of required expression is (1/p)

Regards
Arun
4 years ago
Dear student
 
As we know
 
Cosec²x - cot²x = 1
(Cosecx - cotx) (cosec x + cotx) = 1
 
Hence (cosecx + cotx) = 1 /( cosecx - cotx) = 1/p
 
Hope it helps